# Integrating Neural Operators with Diffusion Models Improves Spectral Representation in Turbulence Modeling

### v1

@@ Abstract
We integrate neural operators with diffusion models to address the spectral limitations of neural operators in surrogate modeling of turbulent flows. While neural operators offer computational efficiency, they exhibit deficiencies in capturing high-frequency flow dynamics, resulting in overly smooth approximations. To overcome this, we condition diffusion models on neural operators to enhance the resolution of turbulent structures. Our approach is validated for different neural operators on diverse datasets, including a high Reynolds number jet flow simulation and experimental Schlieren velocimetry. The proposed method significantly improves the alignment of predicted energy spectra with true distributions compared to neural operators alone. Additionally, proper orthogonal decomposition analysis demonstrates enhanced spectral fidelity in space-time. This work establishes a new paradigm for combining generative models with neural operators to advance surrogate modeling of turbulent systems, and it can be used in other scientific applications that involve microstructure and high-frequency content. See our project page: vivekoommen.github.io/NO_DM
@@ 1 Introduction
Another approach in assimilating data and provide faster solutions than DNS, which is prohibitively expensive at high Reynolds numbers, is through neural operators. In particular, classical DNS provides the solution for a given set of conditions. If there is a change in one of the given conditions, the solvers have to re-run which further adds to the computational cost. To circumvent this issue, neural operators were developed to work with a plurality of operating conditions and even different parameters, e.g. the Reynolds number $Re$ [13, 14, 15, 16, 17, 18]. In a neural operator the objective is to learn the mapping between two or more infinite-dimensional functional spaces. Although training a neural operator offline is expensive, the trained neural operator can infer solutions of unseen conditions almost instantaneously online, with negligible computational costs. Hence, neural operators can serve as efficient surrogates for different applications [19, 20, 21, 22, 23, 16, 24, 25].
Despite the promising results, inference of nonlinear dynamics and its temporal evolution using neural operators suffer from over-smoothing and loss of its high frequency content, a critical property of turbulent flows. This can also be attributed to the spectral bias of neural networks, which is the behavior of neural networks to learn only the low frequencies while remaining oblivious to higher frequency components. The spectral bias of neural operators/networks becomes increasingly significant in the surrogate modeling of turbulent systems whose energy spectrum decays rather slowly leading to non-trivial energy levels at higher wave numbers. In recent literature, the spectral bias has been addressed in [26] who used iterative refinements and [27] who blended neural operators with relaxation methods to alleviate the spectral bias of neural operators. Other ideas implemented for PINNs can be found in [28, 29].
@@
Recently, diffusion models have seen rapid development. Ho et al. [33] introduced the Denoising Diffusion Probabilistic Models (DDPM), which involves corruption of the signal with Gaussian noise followed by sequential denoising for retrieving signal at least from within the same distribution. Although DDPMs achieve high sample quality, the computational cost is high due to its requirement of a large number of sampling steps, typically 1000. Song et al. [34] demonstrated improved training and sampling by using neural networks to accurately estimate the score, which is a denoising-time-dependent gradient field of the perturbed data distribution, and then using numerical stochastic differential equation solvers such as Langevin MCMC (Markov Chain Monte Carlo) to generate samples. Karras et al. [35], performed a systematic investigation that helped elucidate the training and design of diffusion models by analyzing all types of diffusion models from within a common framework. The noise-weighted loss during training and improved sampling routines using preconditioned networks helped to bring down the number of sampling steps from $\approx 1000$ in DDPM to $\approx 30$. Furthermore, [36] demonstrated the efficiency of diffusion models on a variety of image-to-image tasks. For example, in the image colorization problem, a grayscale image serves as a prior for the diffusion model that generates the colorized version. Moreover, in the image inpainting challenge, the masked version of the image serves as an effective prior to the diffusion model that generates the full image. Similarly, [37, 38, 39] used statistically conditioned diffusion models for generating heterogeneous and statistically diverse microstructures. [40] builds stochastic and non-local closure models using conditional diffusion models and neural operators. The works in [41, 42] also present some of the recent applications of diffusion models on scientific data.
@@ 2 Results
Fig.  1: Methodology. First, we train a neural operator to learn mapping between the function spaces by minimizing, typically, the $L^{2}$ norm with respect to available ground truth. Next, we train a diffusion model conditioned on a neural operator’s output to approximate the ground truth data distribution using denoising score matching.
@@ 2.1 Kolmogorov Flow
We train the neural operator ($\mathcal{G}$), to learn the mapping $\omega(x,y,t)_{|t\in[0,10]}\rightarrow\omega(x,y,t)_{|t\in[10,t_{final}]}$ of the vorticity field upto $t=10$ to $t_{final}=20$. We borrow the same problem setup as in [14], but with less smooth initial conditions (see the initial vorticity in Figure 2), causing higher spatial gradients in the resulting simulation as demonstrated in the first row of Figure 2. We investigate the effectiveness of the proposed framework across different neural operator architectures - Fourier Neural Operator (FNO) [14], UNet, and MATCHO [17]. The FNO accurately learns the dominant low wave-number modes, but remains oblivious to the higher wavenumbers corresponding to the finer spatial features. Additionally, the presence of high spatial gradients in the true solution causes the FNO predictions to exhibit Gibbs oscillations. The UNet and MATCHO also show smoothening in their predictions.
The spectral characteristics of the images generated by diffusion models [44, 45] motivated us to utilize them to improve neural operator predictions. Specifically, we train a score-based diffusion model [35] conditioned on the neural operator’s output, to approximate the ground truth data distribution using denoising score-matching. Once trained, the diffusion model uses annealed Langevin dynamics-based sampling to generate its prediction conditioned on the specific neural operator output. Figure 2 illustrates that the vorticity field estimated by the conditioned diffusion model gets rid of the Gibbs oscillations (for FNO), recovers the finer features, and agrees well with the true power spectrum for all the neural operators compared here. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, FNO, and FNO + Diffusion Model in Figure A.1. The proposed framework demonstrates a better alignment with true modes than the neural operator alone.
Fig.  2: Kolmogorov Flow. The first row shows the vorticity field in a kolmogorov flow simulated using a pseudo-spectral solver. The following rows show the predictions of different types of neural operators - FNO, UNet & MATCHO and that of a score-based diffusion model conditioned on the respective neural operator as a prior. The corresponding energy spectra at each timestep are also displayed.
@@ 2.2 Buoyancy-driven Transport
In Figure 4, we visualize the ground truth simulation in the first row, predictions of the neural operator in the second row, and that of the diffusion models conditioned on the neural operator’s output in the third row. We compare the spectral distribution of each of the simulations in row 4. We can observe the over-smoothening in the fields predicted by the neural operator. The diffusion model conditioned on the neural operator achieves better spectral similarity with the ground truth, compared to the neural operator alone. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, UNet, and UNet + Diffusion Model in Figure A.2. Again, the proposed framework demonstrates a better alignment with true modes than the neural operator alone.
@@ 2.3 Turbulent airfoil wake
Here we consider a turbulent wake flow downstream of a NACA0012 airfoil with a chord length of $L_{c}$ at a Reynolds number of 23000, a free stream Mach number of 0.3 and an angle of attack of $6^{\circ}$ [49]. The flow displays coherent structures linked to the Kelvin-Helmholtz instability over the separation bubble and von Karman vortex shedding in the wake, while preserving the broad complexity typical of turbulent flow [50]. This makes it an excellent platform for CFD validation, flow analysis, and the investigation of reduced-complexity models. We directly utilize the large eddy simulation (LES) dataset generated by Towne et al. [50] from "Deep Blue Data", a publicly available database from the University of Michigan. The flow was simulated using a finite-volume based compressible solver - CharLES [51]. The LES dataset is comprised of 16,000 time-resolved snapshots along the mid-span slice, collected at a constant convective time increment of $\tau=0.0104\frac{L_{c}}{U_{\infty}}$. For more details on the problem setup and dataset generation, please refer to section VII in [50].
For training the neural operator we consider a subset of the streamwise velocity component ($u$) of the previously described dataset, consisting of 1000 timesteps, at a time increment of $5\tau$, within a subdomain $4L_{c}\times 2.5L_{c}$. Specifically, we train a U-Net-based neural operator, MATCHO [17], to learn the mapping $u(\bm{x},t)\rightarrow u(\bm{x},t+\Delta t)|_{\Delta t\in[0,50\tau]}$. We split the 1000 timesteps into train, validation and test datasets in the ratio 80:10:10. From Figure 5, we observe that the neural operator predicts a smoothened representation of the true LES simulations. The diffusion model conditioned on the neural operator’s output can recover the finer features, with a better aligned power spectrum. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, MATCHO, and MATCHO + Diffusion Model in Figure A.3. The diffusion model conditioned on the neural operator demonstrates a better alignment with true modes than the neural operator alone.
Fig.  5: Turbulent wake of NACA0012. Distribution of the streamwise velocity $u$ at the mid-span slice at $Re=23,000$. The first row is the ground truth generated from LES, second row is the prediction of neural operator (MATCHO) and the third row is the prediction of a diffusion model conditioned on the neural operator’s output. The last row plots the energy spectrum of the ground truth and predictions.
@@
For training the neural operator we consider the first 1000 timesteps of the velocity field vector ($\bm{u}$) comprising of the axial ($u_{x}$), radial ($u_{r}$) and azimuthal ($u_{\theta}$) components, and utilize the first 800 for training, the next 100 for validation, and the next 100 for testing. The dataset domain is $11.71D$ along the axial direction and $2.78D$ along the radial direction. The neural operator, MATCHO, is trained to learn the mapping $\bm{u}(\bm{x},t)\rightarrow\bm{u}(\bm{x},t+\Delta t)|_{\Delta t\in[0,10\tau]}$.
We illustrate the isosurfaces corresponding to Q-criterion=1.5 in Figure 6. The neural operator is unable to retain the higher wavenumbers in its predictions, especially at the later time states, but captures the general trend of the system contained in the lower wavenumbers. This facilitates the neural operator to behave as an effective prior to the diffusion model which recovers the high frequency components, from angular cross-sections of the velocity vector field inferred by the neural operator. This phenomenon is also illustrated using the power-spectrum plots of the magnitude of the velocity field at different axial cross-sections. Additionally, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, MATCHO, and MATCHO + Diffusion Model in Figure A.4. The diffusion model conditioned on the MATCHO demonstrates a better alignment with true modes than the MATCHO alone.
Fig.  6: 3D turbulent jet. Isosurfaces corresponding to Q-criterion = 1.5 of a turbulent jet ($M=0.9$, $Re=10^{6}$) is visualized. The color of the isosurfaces corresponds to the magnitude of the velocity. The first three rows represent solutions of the reference LES, neural operator (MATCHO), and diffusion model conditioned on the neural operator. The last row compares the energy spectra at different axial cross-sections corresponding to $\Delta t=4\tau,8\tau$.
@@ 2.5 Schlieren velocimetry of turbulent helium jet in air
So far our results demonstrate the effectiveness of using neural operators as priors of diffusion models, on numerically simulated turbulent databases. In the last test case, we investigate if the proposed framework can be directly applied to real experimental setups. We train the surrogate network to model a turbulent helium jet in air from Schlieren velocimetry data [52].
@@
We directly utilize the publically available schlieren velocimetry dataset provided by [52], in raw .tif format. Each snapshot is separated by $\tau=1/6000s$. For training the neural operator we consider 1000 snapshots of the density gradient field ($f$), and utilize the first 800 for training, next 100 for validation and the last 100 for testing. The neural operator, MATCHO, is trained to learn the mapping $f(\bm{x},t)\rightarrow f(\bm{x},t+\Delta t)|_{\Delta t\in[0,10\tau]}$, and the diffusion model takes the output of MATCHO as its prior.
Figure 7 a) illustrates the experimental setup of a traditional single-mirror schlieren apparatus used in [52]. In part b) the first row represents the density gradient captured using the schlieren velocimetry setup at $\Delta t=4\tau,8\tau$, with nozzle on the left and helium expelled towards right along the x-axis. The second and the third row represents the predictions of neural operator and the diffusion model conditioned on neural operator, again at the same time steps. The color bar represents the density gradient normalized between 0 and 1. In the last row we compare the energy spectra of as a function of wavenumber for the three approaches discussed for both the timesteps. From both the contour plots as well as the energy spectrum, the predictions of diffusion model conditioned on the neural operator align better with the ground truth than the neural operator alone. We also perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, MATCHO, and MATCHO + Diffusion Model in Figure 8. The diffusion model conditioned on the neural operator demonstrates a better alignment with true modes than the neural operator alone.
Fig.  7: Schlieren velocimetry. a) Schematic of knife-based single mirror Schlieren apparatus (screenshot from [52]) Settles et al. uses to visualize turbulent jet of helium in the air, with a jet exit $M=0.45$ and $Re=5890$. b) The density gradients at $\Delta t=4\tau,8\tau$ captured using the schlieren velocimetry apparatus (row 1), estimated by neural operator, MATCHO (row 2) and estimated using diffusion model conditioned on neural operator in (row 3). In the last row, we compare the energy spectra of the respective snapshots captured/estimated by the three approaches.

### v2

@@ Abstract
We integrate neural operators with diffusion models to address the spectral limitations of neural operators in surrogate modeling of turbulent flows. While neural operators offer computational efficiency, they exhibit deficiencies in capturing high-frequency flow dynamics, resulting in overly smooth approximations. To overcome this, we condition diffusion models on neural operators to enhance the resolution of turbulent structures. Our approach is validated for different neural operators on diverse datasets, including a high Reynolds number jet flow simulation and experimental Schlieren velocimetry. The proposed method significantly improves the alignment of predicted energy spectra with true distributions compared to neural operators alone. This enables the diffusion models to stabilize longer forecasts through diffusion-corrected autoregressive rollouts, as we demonstrate in this work. Additionally, proper orthogonal decomposition analysis demonstrates enhanced spectral fidelity in space-time. This work establishes a new paradigm for combining generative models with neural operators to advance surrogate modeling of turbulent systems, and it can be used in other scientific applications that involve microstructure and high-frequency content. See our project page: https://vivekoommen.github.io/NO_DM/
@@ 1 Introduction
Another approach in assimilating data and provide faster solutions than DNS, which is prohibitively expensive at high Reynolds numbers, is through neural operators. In particular, classical DNS provides the solution for a given set of conditions. If there is a change in one of the given conditions, the solvers have to re-run which further adds to the computational cost. To circumvent this issue, neural operators were developed to work with a plurality of operating conditions and even different parameters, e.g. the Reynolds number $Re$ [13, 14, 15, 16, 17, 18]. In a neural operator the objective is to learn the mapping between two or more infinite-dimensional functional spaces. Although training a neural operator offline is expensive, the trained neural operator can infer solutions of unseen conditions almost instantaneously online, with negligible computational costs. Hence, neural operators can serve as efficient surrogates for different applications [19, 20, 21, 22, 23, 16, 24, 25, 26].
Despite the promising results, inference of nonlinear dynamics and its temporal evolution using neural operators suffer from over-smoothing and loss of its high frequency content, a critical property of turbulent flows. This can also be attributed to the spectral bias of neural networks [27], which is the behavior of neural networks to learn only the low frequencies while remaining oblivious to higher frequency components. The spectral bias of neural operators/networks becomes increasingly significant in the surrogate modeling of turbulent systems whose energy spectrum decays rather slowly leading to non-trivial energy levels at higher wave numbers. Some of the recent efforts related to mitigating spectral bias suffered by neural networks include multi-scale DNN [28], phase-shift DNN [29] and the use of Fourier feature mappings [30, 31]. In the context of neural operators, PDE-Refiner [32] relies on iterative refinements, [33] blends neural operators with relaxation methods and [34] proposes a novel binned spectral power loss-based training to address the spectral limitations.
@@
Recently, diffusion models have seen rapid development. Ho et al. [38] introduced the Denoising Diffusion Probabilistic Models (DDPM), which involves corruption of the signal with Gaussian noise followed by sequential denoising for retrieving signal at least from within the same distribution. Although DDPMs achieve high sample quality, the computational cost is high due to its requirement of a large number of sampling steps, typically 1000. Song et al. [39] demonstrated improved training and sampling by using neural networks to accurately estimate the score, which is a denoising-time-dependent gradient field of the perturbed data distribution, and then using numerical stochastic differential equation solvers such as Langevin MCMC (Markov Chain Monte Carlo) to generate samples. Karras et al. [40], performed a systematic investigation that helped elucidate the training and design of diffusion models by analyzing all types of diffusion models from within a common framework. The noise-weighted loss during training and improved sampling routines using preconditioned networks helped to bring down the number of sampling steps from $\approx 1000$ in DDPM to $\approx 30$. Furthermore, [41] demonstrated the efficiency of diffusion models on a variety of image-to-image tasks. For example, in the image colorization problem, a grayscale image serves as a prior for the diffusion model that generates the colorized version. Moreover, in the image inpainting challenge, the masked version of the image serves as an effective prior to the diffusion model that generates the full image. Similarly, [42, 43, 44] used statistically conditioned diffusion models for generating heterogeneous and statistically diverse microstructures. [45] builds stochastic and non-local closure models using conditional diffusion models and neural operators. [46, 47, 48, 49, 50] also demonstrate applications of generative AI in scientific modeling.
@@ 2 Results
Fig.  1: Methodology. First, we train a neural operator to learn the mapping between the function spaces by minimizing, typically, the $L^{2}$ norm with respect to available ground truth. Next, we train a diffusion model conditioned on a neural operator’s output to approximate the ground truth data distribution using denoising score matching.
@@ 2.1 Kolmogorov Flow
We train the neural operator ($\mathcal{G}$), to learn the mapping $\omega(x,y,t)_{|t\in[0,10]}\rightarrow\omega(x,y,t)_{|t\in[10,t_{final}]}$ of the vorticity field upto $t=10$ to $t_{final}=20$. We borrow the same problem setup as in [14], but with less smooth initial conditions (see the initial vorticity in Figure 2), causing higher spatial gradients in the resulting simulation as demonstrated in the first row of Figure 2. We investigate the effectiveness of the proposed framework across different neural operator architectures - Fourier Neural Operator (FNO) [14], UNet, and Time-Conditioned UNet (TC-UNet) [17]. The FNO accurately learns the dominant low wave-number modes, but remains oblivious to the higher wavenumbers corresponding to the finer spatial features. Additionally, the presence of high spatial gradients in the true solution causes the FNO predictions to exhibit Gibbs oscillations. The UNet and TC-UNet also show smoothening in their predictions.
The spectral characteristics of the images generated by diffusion models [52, 53] motivated us to utilize them to improve neural operator predictions. Specifically, we train a score-based diffusion model [40] conditioned on the neural operator’s output, to approximate the ground truth data distribution using denoising score-matching. Once trained, the diffusion model uses annealed Langevin dynamics-based sampling to generate its prediction conditioned on the specific neural operator output. Figure 2 illustrates that the vorticity field estimated by the conditioned diffusion model gets rid of the Gibbs oscillations (for FNO), recovers the finer features, and agrees well with the true power spectrum for all the neural operators compared here. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, FNO, and FNO + Diffusion Model in Figure 10. The proposed framework demonstrates a better alignment with true modes than the neural operator alone.
Fig.  2: Kolmogorov Flow. The first row shows the vorticity field in a kolmogorov flow simulated using a pseudo-spectral solver. The following rows show the predictions of different types of neural operators - FNO, UNet & TC-UNet and that of a score-based diffusion model conditioned on the respective neural operator as a prior. The corresponding energy spectra at each timestep are also displayed.
@@ 2.2 Buoyancy-driven Transport
In Figure 4, we visualize the ground truth simulation in the first row, predictions of the neural operator in the second row, and that of the diffusion models conditioned on the neural operator’s output in the third row. We compare the spectral distribution of each of the simulations in row 4. We can observe the over-smoothening in the fields predicted by the neural operator. The diffusion model conditioned on the neural operator achieves better spectral similarity with the ground truth, compared to the neural operator alone. Additionally, we investigate the impact of the neural operator on the diffusion model’s prediction by varying the number of trainable parameters of the neural operator in section 3. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, UNet, and UNet + Diffusion Model in Figure A.1. The proposed framework demonstrates a better alignment with true modes than the neural operator alone. Although the buoyancy-driven flow with $\nu=10^{-2}$ is not an example of a turbulent system, this test case effectively demonstrates the spectral bias suffered by neural operators/networks towards low frequencies.
@@ 2.3 Turbulent airfoil wake
Here we consider a turbulent wake flow downstream of a NACA0012 airfoil with a chord length of $L_{c}$ at a Reynolds number of 23000, a free stream Mach number of 0.3 and an angle of attack of $6^{\circ}$ [57]. The flow displays coherent structures linked to the Kelvin-Helmholtz instability over the separation bubble and von Karman vortex shedding in the wake, while preserving the broad complexity typical of turbulent flow [58]. This makes it an excellent platform for CFD validation, flow analysis, and the investigation of reduced-complexity models. We directly utilize the large eddy simulation (LES) dataset generated by Towne et al. [58] from "Deep Blue Data", a publicly available database from the University of Michigan. The three-dimensional flow was simulated using a finite-volume based compressible solver - CharLES [59]. The LES dataset is comprised of 16,000 time-resolved snapshots collected at a constant convective time increment of $\tau=0.0104\frac{L_{c}}{U_{\infty}}$. For more details on the problem setup and dataset generation, please refer to section VII in [58].
For training the neural operator and the diffusion model we consider only a subset of the streamwise velocity component ($u$) along the mid-span airfoil slice of the aforementioned 3D simulation. We consider 1000 timesteps, at a time increment of $5\tau$, within a subdomain $4L_{c}\times 2.5L_{c}$. We train a UNet-based neural operator, TC-UNet [60], that takes a history of states - $\{u_{t-50\tau},u_{t-45\tau},...,u_{t}\}$ as the input and directly predicts the states - $\{u_{t+5\tau},u_{t+10\tau}\}$. During training, the ground truth from the LES is provided as the input to the neural operator. However, during validation and testing, we carry out an autoregressive rollout by feeding back the states predicted by the neural operator as the input to predict all the states till $t+50\tau$ that corresponds to 80% of the Strouhal time period ($T_{st}$). We split the 1000 timesteps into train, validation and test datasets in the ratio 80:10:10. From Figure 5, we observe that the neural operator predicts a smoothened representation of the true LES simulations. The diffusion model conditioned on the neural operator’s output can recover the finer features, with a better aligned power spectrum. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, TC-UNet, and TC-UNet + Diffusion Model in Figure A.2. The diffusion model conditioned on the neural operator demonstrates a better alignment with true modes than the neural operator alone.
Fig.  5: Turbulent wake of NACA0012. Distribution of the streamwise velocity $u$ at the mid-span slice at $Re=23,000$. The first row is the ground truth generated from LES, second row is the prediction of neural operator (TC-UNet) and the third row is the prediction of a diffusion model conditioned on the neural operator’s output. The last row plots the energy spectrum of the ground truth and predictions.
@@
For training the neural operator we consider the first 1000 timesteps of the velocity field vector ($\bm{u}$) comprising of the axial ($u_{x}$), radial ($u_{r}$) and azimuthal ($u_{\theta}$) components, and utilize the first 800 for training, the next 100 for validation, and the next 100 for testing. The dataset domain is $11.71D$ along the axial direction and $2.78D$ along the radial direction. The neural operator, TC-UNet, is trained to learn the mapping $\bm{u}(\bm{x},t)\rightarrow\bm{u}(\bm{x},t+\Delta t)|_{\Delta t\in[0,10\tau]}$.
We illustrate the isosurfaces corresponding to Q-criterion=1.5 in Figure 6. The neural operator is unable to retain the higher wavenumbers in its predictions, especially at the later time states, but captures the general trend of the system contained in the lower wavenumbers. This facilitates the neural operator to behave as an effective prior to the diffusion model which recovers the high frequency components, from angular cross-sections of the velocity vector field inferred by the neural operator. This phenomenon is also illustrated using the power-spectrum plots of the magnitude of the velocity field at different axial cross-sections. Additionally, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, TC-UNet, and TC-UNet + Diffusion Model in Figure A.3. The diffusion model conditioned on the TC-UNet demonstrates a better alignment with true modes than the TC-UNet alone.
Fig.  6: 3D turbulent jet. Isosurfaces corresponding to Q-criterion = 1.5 of a turbulent jet ($M=0.9$, $Re=10^{6}$) is visualized. The color of the isosurfaces corresponds to the magnitude of the velocity. The first three rows represent solutions of the reference LES, neural operator (TC-UNet), and diffusion model conditioned on the neural operator. The last row compares the energy spectra at different axial cross-sections corresponding to $\Delta t=4\tau,8\tau$.
@@ 2.5 Schlieren velocimetry of turbulent helium jet in air
So far our results demonstrate the effectiveness of using neural operators as priors of diffusion models, on numerically simulated turbulent databases. In the last test case, we investigate if the proposed framework can be directly applied to real experimental setups.
We train the surrogate network to model a turbulent helium jet in air from Schlieren velocimetry data [61].
@@
We directly utilize the publically available schlieren velocimetry dataset provided by [61], in raw .tif format. Each snapshot is separated by $\tau=1/6000s$. For training the neural operator we consider 1000 snapshots of the density gradient field ($f$), and utilize the first 800 for training, next 100 for validation and the last 100 for testing. The neural operator, TC-UNet, is trained to learn the mapping $f(\bm{x},t)\rightarrow f(\bm{x},t+\Delta t)|_{\Delta t\in[0,10\tau]}$, and the diffusion model takes the output of TC-UNet as its prior.
Figure 7 a) illustrates the experimental setup of a traditional single-mirror schlieren apparatus used in [61]. In part b) the first row represents the density gradient captured using the schlieren velocimetry setup at $\Delta t=4\tau,8\tau$, with nozzle on the left and helium expelled towards right along the x-axis. The second and the third row represents the predictions of neural operator and the diffusion model conditioned on neural operator, again at the same time steps. The color bar represents the density gradient normalized between 0 and 1. In the last row we compare the energy spectra of as a function of wavenumber for the three approaches discussed for both the timesteps. From both the contour plots as well as the energy spectrum, the predictions of the diffusion model conditioned on the neural operator align better with the ground truth than the neural operator alone. We also perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, TC-UNet, and TC-UNet + Diffusion Model in Figure A.4. The diffusion model conditioned on the neural operator demonstrates a better alignment with true modes than the neural operator alone.
Fig.  7: Schlieren velocimetry. a) Schematic of knife-based single mirror Schlieren apparatus (screenshot from [61]) Settles et al. uses to visualize turbulent jet of helium in the air, with a jet exit $M=0.45$ and $Re=5890$. b) The density gradients at $\Delta t=4\tau,8\tau$ captured using the schlieren velocimetry apparatus (row 1), estimated by neural operator, TC-UNet (row 2) and estimated using diffusion model conditioned on neural operator in (row 3). In the last row, we compare the energy spectra of the respective snapshots captured/estimated by the three approaches.